Flexible Budget for Selling and Administrative Expenses for a Service Company

Cloud Productivity Inc. uses flexible budgets that are based on the following data:

Sales commissions 15% of sales
Advertising expense 20% of sales
Miscellaneous administrative expense $5,000 per month plus 12% of sales
Office salaries expense $30,000 per month
Customer support expenses $13,000 per month plus 20% of sales
Research and development expense $30,000 per month

Prepare a flexible selling and administrative expenses budget for March for sales volumes of $400,000, $500,000, and $600,000. (Use Exhibit 5 as a model.)

Cloud Productivity Inc.
Flexible Selling and Administrative Expenses Budget
For the Month Ending March 31
Total sales $400000 $500000 $600000
Variable cost:
Sales commissions $ $ $
Advertising expense
Miscellaneous administrative expense
Customer support expenses
Total variable cost $ $ $
Fixed cost:
Miscellaneous administrative expense $ $ $
Office salaries expense
Customer support expenses
Research and development expense
Total fixed cost $ $ $
Total selling and administrative expenses
Answer #1

Answer

  • Requirement

Total sales

$400,000

$500,000

$600,000

Variable cost:

Sales commissions

$60,000

$75,000

$90,000

Advertising expense

$80,000

$100,000

$120,000

Miscellaneous administrative expense

$48,000

$60,000

$72,000

Customer support expenses

$80,000

$100,000

$120,000

Total variable cost

$268,000

$335,000

$402,000

Fixed cost:

Miscellaneous administrative expense

$5,000

$5,000

$5,000

Office salaries expense

$30,000

$30,000

$30,000

Customer support expenses

$13,000

$13,000

$13,000

Research and development expense

$30,000

$30,000

$30,000

Total fixed cost

$78,000

$78,000

$78,000

Total selling and administrative expenses

$346,000

$413,000

$480,000

--Working for Variable cost

Total sales

400000

500000

600000

Variable cost:

Sales commissions

=400000*15%

=500000*15%

=600000*15%

Advertising expense

=400000*20%

=500000*20%

=600000*20%

Miscellaneous administrative expense

=400000*12%

=500000*12%

=600000*12%

Customer support expenses

=400000*20%

=500000*20%

=600000*20%